OpenStudy (anonymous):

Simplify: pi * R^3 + ( - pi * (R-r)^3)

OpenStudy (anonymous):

a^3-b^3=(a-b)(a^2+ab+b^2) use it.

OpenStudy (anonymous):

and take -pi as common?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-pi ((r^3) - (R-r)^3) (R^2 + (R^2-Rr) + (R-r)^2)

OpenStudy (anonymous):

omg how can this be the CENTER OF A MASS?

OpenStudy (anonymous):

this is how i came to this position: "Xcom = (m1*x1 + m2*x2)/ (m1 + m2) Area of Big Circle (m1): m1=Pi* R^2 x1= Center of mass: R m1* x1 = pi * R^3 Therefore value of outer cirle has a full mass of m1 = pi * R^3 -------------------------- Area of Small Circle: (hole, therefore Negative) (m2) m2= -pi * (R-r)^2 x2 = Center of mass: (R-r) m2* x2 = - pi * (R-r)^3 I added up both Areas: The small circle: A, and outer cirle with a small cirle cut out of it.: B, I added up both: (m1*x1 + m2*x2) Total (Mi*Xi) = pi * R^3 + ( - pi * (R-r)^3)" ---- mathematically did i do anything wrong?

OpenStudy (anonymous):

\[\pi \times R ^{3} - \pi \times (R-r) ^{3} = \pi \times (R ^{3}-R^{3}+3R^{2}r+3Rr^{2}-r^{3})\] \[\pi \times (3Rr(R+r)-r^{3})=\pi \times r(3R(R+r)-r^{2})=\pi \times 3R^{2}r+\pi \times 3Rr^{2}-\pi r^{3}\]

OpenStudy (anonymous):

@immeen04 , I think Kathi26 solved it